Belfry Bulletin

Search Our Site

Article Index

Rose Cottage Cave  - Connecting the Entrances and Other News

By Tony Jarratt

“Well, if you know of a better ‘ole, go to it.“  -  C. B. Bairnsfather

Continued from BBs 522-527.

Further Digging: - 15/12/06 – 30/12/06 and the connection.

On 15th December J.C. and T.J. pumped out the pool in the Inlet Tube of the Surface Shaft and filled around eighteen bags with sloppy, gravelly mud. Two days later these two were joined by T.H. and 43 loads were hauled out. Another thirteen bags were filled by T.J. on the 18th and stacked ready for removal. R.W. and T.A. continued with the ginging project. On the freezing cold night of the 20th, after the team had hauled out 15 loads, P.B. and T.J. dug and stacked over a dozen full bags of squalor from the Inlet Tube while H.B, P.C, S.H. and P.H. established further hammering contact from PaulÙs Personal Project in the main cave but were still baffled by the direction in which to dig. The annoying puddle here was drained and 12 skip-loads were hauled out. J.C. filled six bags in the Inlet Tube on the 22nd during a session of “soggy digging” and later that day T.J. filled eleven skips at P.P.P. where he attacked a new site about half a metre above the dig by the drained puddle. This seemed to be a steeply ascending phreatic bedding plane with a stringer of grey organic mud – identical to the Inlet Tube and a good contender for the connection. Masses of heavily calcited breccia blocked the route and were almost impossible to hammer off so next day T.J. and J.C. returned armed with the Hilti drill and detonators. An experimental session of  “micro-blasting” resulted in the removal of two skip-loads of breccia and rock and considerable confusion on the wisdom of wiring detonators in series and using a flat firing battery. The resulting supposed misfire was checked by T.J. and J.C. on Christmas Eve and they were relieved to find that the detonator had actually gone off and was intact enough to provide the latter with an interesting festive pendant. Some token digging was done but the warmth of the Belfry soon lured them away.

With T.J. off to Sutherland for a weekÙs digging at Rana Hole this left J.C. to continue alone at the Surface Shaft side and on the 28th she filled several bags with slop and frustratedly wrote in the log book “How much further can it be?” She was  joined next day by T.H. when the misery of the Inlet Tube drove them to P.P.P. A suggestion that an iron fence rod should be thrust into the dig face just above the micro-blasted site led to J.C. returning to the Inlet Tube where she almost got her eye poked out when the eternally over-enthusiastic T.H. applied this technique! Another hour or so of digging enabled the connection to be made and the first exchange trip done. H.B. and Doug Harris (MCG) did the second one following a drilling session at Plan B Dig on the 30th. This was the last trip of the year.

Halfway and Plan B Digs:- 3/1/07 - 9/5/07.

A major clear up at the Inlet tube and P.P.P. took place on the 3rd January when over 70 loads were hauled out via both entrances and much of the digging gear removed. The tidying team was M.S, H.D, H.B, B.S, Hannah B, S.H, T.J, A.V. and P.H. Halfway Dig was visited by J.B, J.N. and P.C. who continued clearing spoil from the site and gave thanks that the connection dig was over so that they could recruit the staff. R.W continued with the entrance walling on the following day.

Plan B Dig was banged by H.B. and T.H. on the 7th while J.C. and T.J. cleared rocks from Halfway Dig and dumped them further down the cave in any convenient space available. On the following day R.W. continued drystone walling while T.A. extended the length of the steel ladder in preparation for its permanent emplacement in the Surface Shaft entrance. More clearing of both digs took place on the 10th by H.B, H.D, T.J, P.C, B.O, A.L. and Hannah B. On the 14th January T.J. and J.C. continued removing rocks from Halfway Dig and also pumped out the pool in the lower Surface Shaft dig in order to flush out the connection crawl. The Surface Shaft steel ladder was finally bolted and cemented in position by T.A. and R.W. on the 15th. Two days later H.D. and H.B. drilled 13 shot-holes at Plan B Dig and fired a mighty charge while at Halfway Dig J.N, P.C. and A.L. continued clearing while P.B. gave them the benefit of his recently perfected dig engineering knowledge. Four other regulars succeeded in escaping the mayhem to visit other clubsÙ digs – J.C. and T.J. to bang the West Passage chokes in Upper Flood and S.H. and P.H. to assist the Shepton in Gibbets Brow Shaft. More digging and clearing at Halfway took place on the 22nd January when J.B. and T.J. shifted a large amount of mud and rock and measured the spoil rift for future scaffolding. Two short lengths of this were taken down on the 24th by J.B, T.J. and Rob Harper. More clearing was done at Halfway and a three shot-hole cord charge was laid in two boulders to give future working space. H.B. and H.D. continued clearing at Plan B Dig and fired the Halfway charge on their way out. The Henries were back at Plan B on the 28th but due to damage of the electric cable caused by rocks thrown down from Halfway Dig no drilling could be done so they assisted T.H. and T.J. with disposal of said rocks. The bang debris was cleared from Halfway and another charge fired on the half-boulder remaining.

The 29th January saw R.W. and T.A, briefly assisted by T.J, continuing with the shaft walling, spoil heap walling and general tidying up of the site. Two days later H.B. and T.J. fitted the scaffold bars in the Halfway Dig spoil rift. H.B. then drilled thirteen shot-holes at Plan B Dig while T.J. and P.C. continued clearing Halfway. The former then joined H.B. below to assist in charging the holes with mixed cord before firing on the way out. The spoil was cleared on February 11th by H.B. and Rich Bayfield who then “drilled for England”. Three days later H.B. and S.H. laid and fired a 14 shot-hole charge here.

The 17th and 18th February saw T.A. at the Surface Shaft where he “Sneaked in some concrete blocks … while Richard wasnÙt there. (Fixed the foundations for the gate).” He and R.W. continued walling next day and on the 20th R.W. fixed the spoil heap wall and steps.

Halfway Dig was visited by J.B, P.C. and P.H. on the 21st when photos were taken and two scaffold poles inserted as shoring. No digging took place due to lack of dumping space.

On the 26th February H.B. and Hannah B. returned to Plan B Dig (after an abortive trip on the 18th when bad air stopped play) and did some clearing in more breathable conditions. The same area was visited by J.C. on 3rd March when she led a tourist trip for Dominic Gane, Steve and Claire Footitt and Peet Stracey. The problem today was exceptionally wet conditions, which soon drove them back to the Belfry. Allegedly the inflowing stream had been coerced to go down the new entrance by T.H.

R.W. and T.A. continued walling on the 5th in equally unpleasant conditions and on the 7th T.A. bolted on the steel grid gate. Later that day T.J, B.S. and J.C. descended this shaft with intent to dig a side passage but they were defeated by a 3m+ deep pool. A quick trip to Bored of the Rings in the old cave and a grovel through the link gained a view of the pool from the far side. The noise of falling and flowing water made the whole cave feel more “alive”. By the 12th March the weather was much improved and R.W. and T.A. continued walling, briefly assisted by T.J. All admired the colourful addition to the landscape in the form of an ancient and decrepit JCB – severely bogged down in its self-made swamp between the Belfry and the cave and bearing recently added “L” plates. Anyone requiring further information should not ask Ben “Driver of the Year” Selway as a smack in the gob often offends!  In its favour it came in useful as an extra prop in a short documentary film commenced on the 14th March by Alistair Koliasnikoff and Predrag “Pedja” Nikolic of Thames Valley University. H.D, S.H, P.B, P,H, J.C. and T.J. were filmed and interviewed either coming out of the original entrance shaft or hauling skips, four of which, full of rocks for R.W.Ùs wall, reached the surface. Some filming was also done in the HuntersÙ but the underground shots failed due to humidity affecting AlistairÙs camera. H.D. and P.B. checked the state of the drill left at Plan B Dig (it was in good condition despite the recent flooding) then helped J.C. dump six bags of poo, which she had filled in the utterly squalid Halfway Dig. The consensus of opinion tonight was that everyone needed a change of digging scenery!

The next bang at Plan B Dig took place on 21st March when H.D, assisted by M.B. fired a six shot-hole charge while many of the team paid heed to the above and escaped to the relative comfort of Caine Hill Shaft (see separate article). A brief visit was paid on the 26th when R.W, T.A, T.J. and T.H. used 38 bags of clay from this shaft to help infill around the new entrance.

The spoil from the Plan B Dig bang was cleared by H.B. and H.D. on the 28th and seven more shot-holes were drilled.

Another 24 loads of Caine Hill Shaft spoil were dumped on the 2nd April by R.W, T.A, T.J, Paul Weston and Mike Hearn. The seven shot-holes were charged and fired by H.D and H.B. on the 4th and cleared on the 11th. Meanwhile, on the surface Ben S. used his resurrected JCB to improve the Priddy Pot water leat. The Henries returned on the 18th to fire an eleven-hole charge, which they cleared on the 25th before repeating the process. Sometime at the end of April Jake B. continued clearing spoil at the Halfway Dig. On the 9th May the Henries were back at Plan B Dig where they cleared more bang spoil and fired a ten-hole charge.

Landscaping of the spoil heaps has been continued by Rich W, Tony A. and the writer using soil and turf resulting from Ben SelwayÙs JCB operations.

To be continued in a future BB.

New and Resurrected Diggers

Doug Harris (MCG), Bob Smith, Hannah Bell, Rob Harper, Rich Bayfield, Matt Blount, Paul Weston, Mike Hearn.

The Old Brigade

Jane Clarke, Tony Jarratt, Trevor Hughes, Pete Hellier, Paul Brock, Phil Coles, Sean Howe, Henry Bennett, Tony Audsley, Rich Witcombe, Martin Smith (OSCG), Henry Dawson, John Noble, Anne Vanderplank, Jake Baynes, Ben Ogbourne, Alex Livingston, Ben Selway.

The Film Team

Alistair Koliasnikoff, Predrag “Pedja” Nikolic, Mario Michalidis, ( Thames Valley University).   


Charlie Adcock (Event Horizon Pyrotechnics).